### v2

# Learning Exposure Correction in Dynamic Scenes

Abstract
Capturing videos with wrong exposure usually produces unsatisfactory visual effects. While image exposure correction is a popular topic, the video counterpart is less explored in the literature. Directly applying prior image-based methods to input videos often results in temporal incoherence with low visual quality. Existing research in this area is also limited by the lack of high-quality benchmark datasets. To address these issues, we construct the first real-world paired video dataset, including both underexposure and overexposure dynamic scenes. To achieve spatial alignment, we utilize two DSLR cameras and a beam splitter to simultaneously capture improper and normal exposure videos. In addition, we propose a Video Exposure Correction Network (VECNet) based on Retinex theory, which incorporates a two-stream illumination learning mechanism to enhance the overexposure and underexposure factors, respectively. The estimated multi-frame reflectance and dual-path illumination components are fused at both feature and image levels, leading to visually appealing results. Experimental results demonstrate that the proposed method outperforms existing image exposure correction and underexposed video enhancement methods. The code and dataset will be available soon.
1 Introduction
Figure 1: The upper part is the benchmark for our proposed video exposure correction. Taking underexposed video as an example, the lower part is the visual comparison with image-based method. LACT [2] takes single frames as input and results in temporal exposure incoherence with low visual quality. Our method utilizes temporal information to maintain the same exposure.
When taking photos or videos across various scenarios and lighting conditions, the results may not always be ideal due to low light, backlight, or intense illumination factors. Images and videos captured with underexposure or overexposure often yield unsatisfactory visual effects. When enhancing such mixed multiple exposure data, there are large discrepancies in the training exposure correction process of the network. This challenge is further compounded when dealing with videos captured in dynamic scenes, where cameras or objects undergo substantial movement. This paper focuses on correcting improperly exposed videos taken from dynamic scenes with suboptimal lighting conditions, which has not been well explored.
In recent years, the development of paired image datasets [3, 1] has spurred the creation of several methods [12, 11, 29, 23, 34] focused on learning image exposure correction through supervised learning. They learn a mapping relation from images taken in underexposure or overexposure conditions to those with normal exposures. However, these methods generally do not handle videos of dynamic scenes, as directly applying them in a frame-by-frame manner often results in temporal incoherence with low visual quality [19, 15], as shown in Fig. 1. A notable portion of some methods [31, 25, 33, 8, 4] is rooted in the Retinex theory [14] to achieve underexposed image or video enhancement. The objective involves modeling the relationship between the illumination and reflectance components through extensive data-driven learning. This cannot be directly applied to multiple exposure correction tasks with opposite optimization goals. To date, there is no current work exploring real-world video exposure correction datasets and methods. Therefore, we propose a benchmark for training and evaluating video multi-exposure correction.
On the one hand, we aim to collect a high-quality video dataset with both underexposure and overexposure, and the main difficulties are as follows. First, collecting paired video datasets requires capturing two videos, one under abnormal exposure conditions and another under normal light of the same dynamic scene and camera motion. Second, it requires accurate alignment of each pair of corresponding frames in the two videos in both spatial and temporal dimensions. To overcome the above challenges, we design a two-camera system with a beam splitter to ensure no parallax between the two cameras with different aperture and ISO parameter settings. Two DSLR cameras simultaneously capture improper and normal exposure videos. Then we perform precise alignment on the captured pairs to make them aligned with each other. Furthermore, the well-exposed videos are manually rendered by professional photographers and then serve as the ground truth of the dataset. Finally, we construct a new dataset of 119 high-quality spatially-aligned videos named as DIME dataset, standing for “Dynamic scenes In Multiple Exposure”. It contains diverse real-world scenes, large camera and object motions, and accurate spatial alignment.
On the other hand, there are varying degrees of improper exposure between videos or between frames. We propose an end-to-end Video Exposure Correction Network (VECNet) to address both overexposure and underexposure problems in videos. Specifically, based on Retinex theory, we introduce the dual-stream illumination mechanism to construct illumination maps and enhanced images for the overexposure and underexposure branches. Then, to learn aligned reflectance components from consecutive frames, we introduce a novel multi-frame alignment strategy. Finally, a two-stage process is proposed to synthesize dual illumination and reflectance for restoration at feature and image levels. Our contributions are as follows:
  • We construct the first high-quality paired video exposure correction dataset for dynamic real-world scenes with multiple exposures, camera and object motions, and precise spatial alignment.
  • We propose an effective exposure correction network based on Retinex theory to enhance overexposed and underexposed videos.
  • We conduct extensive experiments to demonstrate the superiority of our dataset and method.
2 Our DIME Dataset
Figure 2: We built the optical system to capture the over-/under- and normal exposure video pairs.
Figure 3: Statistics on our DIME dataset. (a) is the input-ground truth brightness mapping curve statistics. (b) Luminance distribution for overexposure videos. (c) Luminance distribution for underexposure videos. (d) Luminance distribution for ground truth videos.
To learn video exposure correction tasks by training our model, we need a large number of videos including realistic overexposure and underexposure errors and corresponding paired and properly exposed ground truth videos. As discussed in Sec. 2, such datasets are currently not publicly available to support video exposure correction research. Therefore, our first task is to create a novel dataset that captures various videos in dynamic and real-world scenes.
2.1 Hardware Design
Capturing multiple exposure image pairs for static scenes can be easily realized with a single exposure adjustable DSLR camera. However, For dynamic exposed video capturing, we need to utilize two cameras with different aperture and ISO parameter settings. To keep the cameras in sync, we use an infrared remote control to signal two camera-mounted receivers for simultaneous capture. To prevent parallax problems caused by different shooting positions and ensure the same motion trajectory, inspired by  [9, 25], we use a beam splitter to split the incident light into two beams with a ratio of 1:1 and enter the two camera lenses respectively, as shown in Fig. 2. In order to receive the natural incident light from the same viewpoint, we design and print an opaque 3D model box, which is also used to hold the beam splitter and connect the cameras. These components are concentrated on an optical breadboard and then fixed on the tripod to improve stability.
Figure 4: LOE and optical flow of different datasets.
2.2 Data Collection
We use two Canon EOS R10 cameras to collect paired videos. Specifically,to minimize the loss of detail in shadows and highlights, we set one camera to automatic exposure mode and capture video in Canon Log [^1] https://cam.start.canon/en/C004/manual/html/UG-03_Shooting-2_0080.html format, which retains more visual information throughout the dynamic range. We then hand it over to professional photographers using DaVinci Resolve Studio [^2] https://www.blackmagicdesign.com/products/davinciresolve/studio software for manual rendering, ultimately producing high-quality and normal-exposure videos. The other camera captures low-quality abnormal exposure videos with the commonly used sRGB format. All the other settings are set to default values to simulate real capture scenarios. For each scene, we collect paired short 10-second videos at 50 FPS to capture both slow and fast motions.
Figure 5: Example videos from our DIME dataset cover under-/over-exposures, indoor/outdoor, and camera/object motions.
2.3 Data Processing
Due to the slight errors in manually assembling the above hardware, there is still misalignment between the paired videos. Therefore, we utilize a two-stage frame alignment strategy to obtain aligned pairs. First, we estimate a homography matrix between the under-/overexposed frame and the corresponding normal exposed frame using the matched SIFT [18] key points. In this way, we can roughly crop matching regions in corresponding frames. Then we utilize a traditional flow estimation algorithm called DeepFlow [32] to perform pixel-wise alignment. Finally, we use center cropping to remove alignment artifacts around boundaries, producing precise spatial aligned frames. It should be noted that we only perform alignment correction on normal exposed frames and make the low-quality input of our network consistent with real captured abnormal exposed frames. In addition, we introduce lightness-order-error (LOE) [26] and optical flow [6] to quantitatively evaluate alignment performance and motion activity. As shown in Fig. 4, the paired frames of our dataset are better aligned and have faster movement than that of the low-light enhancement datasets including DRV [5] and SDSD [25], indicating stronger performance in real dynamic scenes.
We totally captured 119 groups of videos with a total of 21,951 frames to form our DIME dataset. Each video consists of 100-200 frames, and the resolution is 1,920 $\times$ 1,080. Fig. 3 presents some statistical analysis results. Fig. 5 gives three examples of our aligned video pairs under under-/over-exposure and normal conditions. Our captured videos vary from indoor to outdoor in real scenes and include camera and object motion types. See more detailed information in the supplementary file.
Figure 6: Overview of our framework. It contains three modules including Multi-frame Phase Alignment (MPA), Dual-stream Illumination Construction (DIC), and Two-stage Synthesis Restoration (TSR).
Figure 7: The detail of the Multi-frame Phase Alignment (MPA) module. We use discrete Fourier transformation to map the frames from pixel space to Fourier space, then adopt the phase branch for alignment.
3 Methodology
3.1 Overview
The Video Exposure Correction (VEC) task can be formulated as seeking a mapping function $\mathcal{F}$, which maps consecutive 8-bit per channel sRGB frames $\mathcal{I}$ to enhanced frames $\mathcal{O}$ such that $\mathcal{O}=\mathcal{F}(\mathcal{I})$. During the training step, given 2N+1 consecutive frames $\mathcal{I}_{[t-N:t+N]}$ captured under different exposure conditions, the center frame is denoted as the reference frame while the others are supporting frames. The proposed end-to-end VECNet aims to restore the exposure of the reference frame $\mathcal{I}_{t}$ with the help of supporting frames, thereby generating the output frame $\mathcal{O}_{t}$. Based on Retinex theory [14], the overall architecture of VECNet is shown in Fig. 6, which is composed of three sub-networks: Multi-frame Phase Alignment Module (MPA), Dual-stream Illumination Construction Unit (DIC) and Two-stage Synthesis Restoration Unit (TSR). MPA module first aligns supporting frames to the reference frame to maintain temporal consistency and then learns the reflectance map together. Next, DIC unit estimates dual-path illumination maps for the reference frame to adaptively adjust its underexposure and overexposure factors. Finally, TSR unit fuses the above illumination and reflectance maps at feature and image levels, obtaining the high-quality output frame $\mathcal{O}_{t}$ with proper exposure.
3.2 Multi-frame Phase Alignment
Directly applying image exposure correction [1, 11] to each frame can cause flickering [25] due to ignoring temporal information. To take advantage of it, we take consecutive frames as the model’s input. However, since the texture information of different frames is misaligned, a direct fusion of multiple frames could cause blur and artifacts [22, 25]. Existing works mainly warp supporting frames to the reference frame by pyramid cascading and deformable convolution [28, 35] or group shift operations [16] directly at the feature level. However, in the video exposure correction task, There are differences in exposure between frames as camera motion produces different light positions. Such exposure mutation interferes with alignment. Recently,  [11] conducts exposure correction from a Fourier-based perspective, i.e., the amplitude component of an image reflects the lightness representation, while the phase component corresponds to structures and is less related to lightness. To this end, we take multi-frame alignment at the phase level in this section, as shown in Fig. 7.
Specifically, given a channel in single frame $\mathit{x}\in\mathbb{R}^{H\times W}$, the Fourier transform $\mathcal{F}$ converts $\mathit{x}$ to a complex component $\mathit{X}$ in Fourier space, which can be written as:
$\begin{split}&\mathcal{F}(x)=\frac{1}{\sqrt{HW}}\sum_{h=0}^{H-1}\sum_{w=0}^{W-1}x(h,w)e^{-j2\pi(\frac{h}{H}u+\frac{w}{W}v)},\end{split}$ (1)
Then, amplitude $\mathcal{A}(\mathit{X})$ and phase $\mathcal{P}(\mathit{X})$ components can be divided from the complex component $\mathit{X}$:
[…]
where $\mathit{R}$ and $\mathit{I}$ denote the real and imaginary parts of the complex $\mathit{X}$. We also apply the same strategy to each channel of input frames and then get consecutive amplitude components $\mathcal{A}(\mathit{X}_{[t-N:t+N]})$ and phase components $\mathcal{P}(\mathit{X}_{[t-N:t+N]})$. Since phase components respond more to the structure information and are less affected by the lightness while amplitude components are just the opposite, the phase components from supporting frames are warped to that of the reference frame by deformable convolution network (DCN). We compute the offset $\delta p^{k}_{t+i},i\in[-N,N],i\neq 0$ learned from corresponding $\mathcal{P}(\mathit{X}_{t+i})$ and $\mathcal{P}(\mathit{X}_{t})$ for the DCN. The aligned phase component $\tilde{\mathcal{P}}(\mathit{X}_{t})$ is obtained with the learned offset. The process can be formulated as:
$\begin{split}\delta p^{k}_{t+i}&=\mathcal{M}_{i}(\mathcal{P}(\mathit{X}_{t+i})\odot\mathcal{P}(\mathit{X}_{t})),\\ \tilde{\mathcal{P}}(\mathit{X}_{t+i})&=\mathcal{G}_{i}(DCN(\mathcal{P}(\mathit{X}_{t+i}),\delta p^{k}_{t+i})),\end{split}$ (3)
where $\odot$ denotes channel concatenation, $\mathcal{M}_{i}$ and $\mathcal{G}_{i}$ are the mapping function composed of several convolution layers. In addition, we obtain $\bar{\mathcal{A}}(X)$ aggregate multi-frame amplitude component to achieve unified representation. Finally, we recombine the aligned phase component and aggregated amplitude component by inverse Fourier transform $\mathcal{F}^{-1}$. The process is formulated as:
$\begin{split}\bar{\mathcal{A}}(X)&=\mathcal{H}[\mathcal{A}(\mathit{X}_{t-N})\odot...\mathcal{A}(\mathit{X}_{t+N})],\\ \tilde{x}_{t+i}&=\mathcal{F}^{-1}(\bar{\mathcal{A}}(X),\tilde{\mathcal{P}}(\mathit{X}_{t+i})),\end{split}$ (4)
where $\mathcal{H}$ is the aggregating function composed of several convolution and relu function layers. Then the aligned features consisting of multiple single-channel $\tilde{x}_{t+i}$ are sent to learn the reflectance map.
3.3 Dual-stream Illumination Construction
To learn adaptive representations for both underexposure and overexposure, we exploit dual-stream illumination construction in our model. From the Retinex theory, the reference frame $\mathcal{I}_{t}$ can be typically decomposed into an illumination map $L$ and a reflectance map $\mathcal{R}$. However, this strategy can only be learned from underexposure conditions as $L$ falls into the values within [0,1]. Existing Retinex-based methods [4, 7] would increase the exposure of the overexposed inputs.
To suppress overexposure while enhancing underexposure, we propose a dual-stream mechanism to achieve illumination construction, sharing parameters with each other. Specifically, as shown in Fig. 6, we treat the antithetic exposure $\mathcal{I}^{inv}_{t}$ as the reverse frame of $\mathcal{I}_{t}$. We first encode $\{\mathcal{I}_{t},\mathcal{I}^{inv}_{t}\}$ to $z_{t},z^{inv}_{t}$ in the latent space. Then we use a U-Net [21] to learn the mapping from $\{z_{t},z^{inv}_{t}\}$ to $\{\tilde{z}_{t},\tilde{z}^{inv}_{t}\}$. Finally we decode $\{\tilde{z}_{t},\tilde{z}^{inv}_{t}\}$ to $\{L_{t},L^{inv}_{t}\}$ in the image space. By calculating with the learned and extended illumination map $L^{inv}$, we can obtain the enhanced image of overexposure $\mathcal{I}^{o}_{t}$. This process can be formulated as:
$\begin{split}\mathcal{I}^{u}_{t}&=\mathcal{I}_{t}\cdot L_{t}^{-1},\\ \mathcal{I}^{inv}_{t}&=1-\mathcal{I}_{t},\\ \mathcal{I}^{o}_{t}&=1-\mathcal{I}^{inv}_{t}\cdot(L_{t}^{inv})^{-1},\end{split}$ (5)
where $\mathcal{I}^{u}_{t}$ represents the enhanced frame of underexposure, $\cdot$ denotes element-wise multiplication.
3.4 Two-stage Synthesis Restoration
[…]
3.5 Objectives
We incorporate a range of loss terms to facilitate the training of our model. We employ the commonly utilized pixel-wise charbonnier loss term, denoted as $\mathcal{L}_{pix}$, to measure the accuracy of pixel reconstruction. Based on Retinex theory, we present reflectance consistency loss $\mathcal{L}_{R}$ between paired generated frames and ground truths. To guide the network to estimate and smooth the dual illumination maps, we apply the total variation [24] loss term $\mathcal{L}_{tv}$. They can be formulated as:
$\begin{split}\mathcal{L}_{R}&=||\mathcal{T}(\mathcal{O}_{t})-\mathcal{T}(\mathcal{I}_{t}^{gt})||^{2}_{F},\\ \mathcal{L}_{tv}&=\sum_{c}[{(\partial_{x/y}L_{t})}^{2}+{(\partial_{{x/y}}L_{t}^{inv})}^{2}],\end{split}$ (7)
where all channels (c) of all pixels are summed, $\partial_{x/y}$ are partial derivatives in horizontal and vertical directions in the image space. In addition, In order to ensure exposure continuity between video frames, we introduce amplitude consistency loss term $\mathcal{L}_{amp}$. It is formulated as:
$\begin{split}\mathcal{L}_{amp}&=||\mathcal{A}(X_{\mathcal{O}_{t}})-\mathcal{A}(X_{I^{gt}_{t}})||_{F}^{1},\end{split}$ (8)
where $X_{(\cdot)}$ represents the output or ground truth normal exposure image in Fourier space. The overall function can be written as:
$\begin{split}\mathcal{L}_{total}&=\lambda_{1}\mathcal{L}_{pix}+\lambda_{2}\mathcal{L}_{R}+\lambda_{3}\mathcal{L}_{tv}+\lambda_{4}\mathcal{L}_{amp},\end{split}$ (9)
where $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$, and $\lambda_{4}$ are weight hyperparameters.
4 Experiments
Figure 8: Qualitative comparisons of the multi-frame results with state-of-the-art methods on the DIME dataset.
4.1 Implementation Details
We implement our framework with Pytorch on a single NVIDIA GeForce RTX 3090 GPU. We use $N=2$ consecutive frames without intervals as input. We divide the DIME dataset into training/validation/test sets according to the number of videos in 90:9:20. Then we use the training set to train our model with a total of 400,000 iters. The parameters of the network are optimized by the ADAM optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.99$ The learning rate is 0.001 while the batch size is 8. The frames apply random horizontal and vertical flipping to augment the input data. We train on patches of size $256\times 256$ pixels. The weights for the terms in the loss function in Eq. (9) are $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}=1.0,1.0,0.01,100.0$. The additional experimental results are presented in the supplementary file.
4.2 Baselines
Since there is currently no related method for video exposure correction, we adopt several image exposure correction methods for comparison including MSEC [1], DRBN-ENC [10], ECLNet [12], FECNet [11], and LACT [2]. We also take low-level tasks adjacent to video exposure correction, including several methods in the fields of video low-light enhancement and video restoration, such as SMOID [13], SDSD [25], RVRT [17], and DIDNet [8].
Method Under Over Average
PSNR SSIM PSNR SSIM PSNR SSIM
SMOID [13] 7.18 0.4633 7.59 0.4315 7.38 0.4452
SDSDNet [25] 14.19 0.6491 18.72 0.8420 16.45 0.7385
RVRT [17] 16.62 0.7859 16.93 0.8555 16.78 0.8207
DIDNet [8] 16.20 0.8435 18.61 0.9022 17.41 0.8729
MSEC [1] 13.22 0.8071 18.01 0.8854 15.61 0.8462
DRBN-ENC [10] 15.69 0.8380 18.02 0.9203 16.83 0.8782
ECLNet [12] 13.16 0.7662 20.77 0.9105 16.88 0.8367
FECNet [11] 14.65 0.8461 22.52 0.9009 18.49 0.8729
LACT [2] 12.91 0.7688 17.62 0.8236 15.26 0.7962
VECNet (Ours) 18.63 0.8987 21.91 0.9379 20.27 0.9183
Table 1: Quantitative PSNR and SSIM results of different methods on Underexposed and Overexposed samples of the DIME dataset. A higher metric indicates better performance. The best and second results are marked in bold and underlined, respectively.
Method DIME LLIV-Phone
NIQE ALV NIQE ALV
SMOID [13] 5.43 35.32 6.76 6.06
SDSDNet [25] 4.35 18.04 4.07 8.99
RVRT [17] 4.17 20.81 5.75 7.07
DIDNet [8] 7.25 42.60 6.93 10.32
MSEC [1] 5.38 37.28 5.80 232.8
DRBN-ENC [10] 4.48 45.51 5.24 129.4
ECLNet [12] 4.40 23.31 5.38 141.0
FECNet [11] 4.52 29.11 5.26 105.5
LACT [2] 5.01 35.57 5.62 171.6
VECNet (Ours) 4.10 15.72 4.01 4.65
Table 2: Quantitative results of different methods in terms of NIQE and ALV. A lower metric indicates better performance. The best and second results are marked in bold and underlined, respectively.
4.3 Quantitative Evaluation
We demonstrate the superiority of our method and the impact of the DIME dataset through experiments in this section. We conduct quantitative experiments on the test set, which consists of 10 underexposed video pairs and 10 overexposed video pairs with a total of 3,824 frames, as described in Sec. 2. We adopt the following three standard metrics to evaluate the pixel-wise accuracy and the perceptual quality of our results: reference-based (i) Peak Signal-to-Noise Ratio (PSNR) and (ii) Structural Similarity Index Measure (SSIM) [30], non-reference based (iii) Natural Image Quality Evaluator (NIQE) [20] and (iv) Average Luminance Variance (ALV) [15] metrics. Table 1 reports the quantitative evaluation results of different methods. As exhibited in the table, our proposed VECNet outperforms all other methods in PSNR and SSIM metrics, demonstrating its superior performance in video exposure correction tasks.
Validation on unpaired videos. To evaluate the generalization ability of our method under challenging scenes with large-scale motions, we also conduct extensive experiments on the LLIV-Phone dataset [15]. Since there is no ground truth, all compared baselines are only tested on the dataset. Note that we report the results of our model trained on the DIME training set without further tuning or re-training on any of these datasets. Table 2 reports the quantitative results. Our method achieves the best results in terms of NIQE and ALV, demonstrating the stability when handling real-world videos with various motions.
Various kinds of device video enhancement. We collect a total of 30 abnormal exposure videos from three devices, 10 from each device. We test them with the compared methods pretrained on the DIME training set. As described in Table 5, our method achieves better results, displaying the effectiveness for more applications. We present the luminance curves of the enhanced video frames to verify the degree and continuity of the exposure, as shown in Fig. 11.
4.4 Qualitative Evaluation
We perform thorough qualitative evaluations on the DIME and LLIV-Phone datasets to assess the performance of our proposed method. We present single-frame and multi-frame results in Fig. 8 and Fig. 9, which indicates that VECNet delivers more natural and reasonable enhancement. In particular, low-light video enhancement methods mainly produce frames with unpleasing regions, leading to substantial loss of texture. Image-based exposure correction methods result in significant color deviation, which adversely affects the visual quality of the images. SMOID and RVRT produce frames with severe blocking artifacts. Moreover, image-based methods generate exposure inconsistent frames as the temporal information has not been well utilized to avoid flickering. After a comprehensive evaluation of the comparative results of different methods on two datasets, our proposed method exhibits excellent visual performance in terms of global brightness, color recovery, and detail while maintaining temporal stability without flickering artifacts and motion blur.
Model MPA DIC TSR PSNR SSIM
phase single dual stage-1 stage-2
(a) 18.20 0.7251
(b) 18.78 0.8728
(c) 19.12 0.9025
(d) 20.27 0.9183
Table 3: Ablation study for investigating the components of the specific modules. For the MPA module, we ablate the phase operation. For the DIC module, we ablate the single and dual streams.
Figure 9: Qualitative comparisons of the multi-frame results with state-of-the-art methods on the LLIV-Phone dataset.
Method Other Same VECNet
SMOID [13] 0.20 0.15 0.65
SDSDNet [25] 0.25 0.20 0.55
RVRT [17] 0.05 0.15 0.80
DIDNet [8] 0.25 0.10 0.65
Table 4: User study score comparisons with baselines. “VECNet” is the percentage that our result is preferred, “Other” is the percentage that some other approach is preferred, “Same” is the percentage that the users have no preference.
Method iPhone XiaoMi Internet Average
LACT [2] 4.42/38.32 4.85/47.50 4.11/24.13 4.46/36.65
DIDNet [8] 4.10/30.90 4.55/27.66 3.72/18.72 4.12/25.76
VECNet 3.44/19.01 4.49/25.83 3.49/8.37 3.81/17.74
Table 5: Quantitative results of different methods for multi-exposure videos captured from various devices and internet, evaluated by NIQE and ALV metrics.
4.5 User Study
We conduct a user study with 30 participants to evaluate the subjective perception of different methods. We select 20 testing videos from the DIME dataset. The videos are then enhanced using five video enhancement methods (SMOID, SDSDNet, RVRT, DIDNet, and VECNet) to perform pairwise comparisons. For a fair comparison, the user study is conducted in the same environment (room, display, and light). Participants are asked to choose: “baseline is better”, “VECNet is better”, or “no preference” in terms of brightness, naturalness, contrast, color, and temporal stability. In Table 4, our method is preferred by human subjects, suggesting that our results are more natural and realistic.
4.6 Ablation Study
We provide a series of ablation studies to evaluate the effectiveness of each component in the proposed method. Note that all experiments are conducted on the DIME dataset. The ablation results are shown in Table 3 and Fig. 10.
For the MPA module, we remove it and directly fuse the original unaligned multiple frames to learn reflection maps to generate the final result. It can be seen MPA brings expected quantitative performance improvement. The results of Fig. 10 (b) are more visually appealing with fewer noises and motion blurs than (a), which further demonstrates the effectiveness of MPA. For the DIC unit, we replace the symmetrical exposure stream with a single path setting. It can be seen the overexposure correction shows little improvement from Fig. 10 (b). For the TSR unit, we split the TSR unit into two separate stages. Comparing Fig. 10 (c) and (d), we can see that our two-stage strategy helps preserve more color and texture information.
Figure 10: Visualizations of ablation study on proposed module.
Figure 11: Lux curves of the video enhanced by different methods. Left is underexposure cases while right is overexposure cases.
5 Conclusion
We build the first high-quality paired video exposure correction dataset for dynamic real-world scenes with multi-exposures, camera and object motions, and precise spatial alignment. Based on the dataset, we propose a method by dealing with the underexposed and overexposed inputs in a dual-stream manner. By utilizing phase alignment and synthesis modules, the exposure of the videos is well corrected and restored. Experiments demonstrate that the proposed method outperforms several image/video methods in low-level tasks adjacent to video exposure correction.